OpenStudy (anonymous):

if f(x)= 4/(x-2), find f '(4)

OpenStudy (dumbcow):

think of it as 4(x-2)^-1, now differentiate using power rule -4(x-2)^-2

OpenStudy (dumbcow):

technically chain rule as well but d/dx(x-2) = 1

OpenStudy (anonymous):

so 1 would be my answer?

OpenStudy (dumbcow):

f'(x) = -4/(x-2)^2 f'(4) means you replace x with 4 =-4/(4-2)^2 =-4/4 = -1

OpenStudy (anonymous):

oh, ok. Thanks. for this function, how would you calculate the slope of the tangent line at the point (4,2)

OpenStudy (dumbcow):

guess what you already did. f'(4) is the slope the tangent line at (4,2)

OpenStudy (anonymous):

crap, not slope. I meant intercept.